OpenStudy (anonymous):

The measure of PQR is 60°. The length of radius QR is 8 inches. What is the area of sector PQR? (Use 3.14 for .)

OpenStudy (campbell_st):

the area is simply the fraction of the circle multiplied by the area of the circle \[A = \frac{x}{360} \times \pi r^2\] x is the sector angle, r = radius... use your information to calculate the answer